x+y=1
2x-y=2
rules say to use substitution to solve the system

Answer by sdmmadam@yahoo.com(530)

x+y=1----(1)
2x-y=2 ----(2)
(1) implies x = (1-y) ----(*)(subtracting y from both the sides)
Putting (*) in (2)
2x-y=2
2(1-y)-y=2
2-2y-y = 2
2-3y = 2
2-2 = 3y
0=3y
That is 3y = 0
y =0/3 = 0
Putting y=0 in (*)
x=(1-y) = (1-0)=1
Answer: x=1,y=0
Verification: Putting x=1 and y=0 in (1)
x+y =1
LHS = x+y = 1+0 = 1 =RHS